determine the number of milliliters of a 70% sulfuric acid solution that must be added to 160 milliliters of a 25% solution to make a 30% solution
Answer by stanbon(75887)

determine the number of milliliters of a 70% sulfuric acid solution that must be added to 160 milliliters of a 25% solution to make a 30% solution
-----
Equation:
acid + acid = acid
0.25*160 + 0.70x = 0.30(160+x)
------
25*160 + 70x = 30*160 + 30x
40x = 5*160
x = 20 ml (amt. of 70% solution to add)
-----------
Cheers,
Stan H.
-------